KrazyKenzie:

A horse has a momentum of 25 kg*m/s and a velocity of 2.5 m/s. What is the horse's mass? a 62.k kg b 10 kg c 10 m/s d 0.1kg

jkmiec26:

What are you thinking?

ZombieLover:

v = velocity. Hence, mass is: m = p / v. From the question: p = 25 kgm/s. v = 2.5 m/s. Mass is: m = 25 / 2.5

jhonyy9:

the momentum ,,p" formula is p = m*v m - the mass v - velocity - from this formula you need calcule the m value p = m*v divide both sides by v and will get m = p/v like above wrote @ZombieLover @Laylalyssa